\begin{thebibliography}{99}


\bibitem{bass}
H. Bass, \emph{On the ubiquity of Gorenstein rings}, Math. Z. \textbf{82} (1963) 8-28.

\bibitem{ca} G.
Casnati, \emph{
Isomorphism types of Artinian Gorenstein local algebras of multiplicity at most 9},
Comm. Algebra \textbf{38} (2010),   2738-2761.



\bibitem{ev}
J. Elias, G. Valla, \emph{
Structure theorems for certain Gorenstein ideals},
Michigan Math. J. \textbf{57} (2008), 269-–292.

\bibitem{ga}
V. Gasharov, I. Peeva, \emph{Boundedness versus periodicity over commutative local rings}, Trans.
Amer. Math. Soc. \textbf{320} (1990), 569-–580.




\bibitem{sga2}
A. Grothendieck,
\emph{Cohomologie locale des faisceaux coh\'{e}rents et th\'{e}orèmes de Lefschetz locaux et globaux} (SGA 2).
S\'{e}minaire de G\'{e}om\'{e}trie Alg\'{e}brique du Bois Marie  1962. Amsterdam: North Holland Pub. Co. (1968).

\bibitem{dil}
T.K.
Gunston, \emph{Cohomological degrees, Dilworth numbers and linear resolution}, Thesis (Ph.D.) Rutgers, Available at arXiv:1008.3711.
\bibitem{wja}T.
	Harima, J.  Migliore, U.  Nagel, J. Watanabe,
\emph{The weak and strong Lefschetz properties for Artinian K-algebras},
J. Algebra \textbf{262} (2003), no. 1, 99--126.

\bibitem{wat3}T.
Harima, T. Maeno, H. Morita, Y. Numata, A.  Wachi,J. Watanabe,
\emph{The Lefschetz properties},
Lecture Notes in Mathematics, \textbf{2080} Springer, Heidelberg, (2013).


\bibitem{iw}
H. Ikeda, J. Watanabe, \emph{The Dilworth lattice of Artinian rings}, J. Commut. Algebra \textbf{1} (2009),  315–-326.

\bibitem{i}
H. Ikeda, \emph{Results on Dilworth and Rees numbers of Artinian local rings}, Japan J. Math. (N.S.) \textbf{22} (1996),  147–-158.

\bibitem{ii}
H. Ikeda, \emph{The Dilworth number and Rees number of one-dimensional modules}, Comm. Algebra \textbf{25} (1997), no. 8, 2627–-2633.



 \bibitem{M}
H. Matsumura, \emph{Some new results on numbers of generators of ideals in local rings},  Topics in algebra, Part 2 (Warsaw, 1988), 157-–161,
Banach Center Publ., 26, Part 2, PWN, Warsaw, (1990).

\bibitem{Mat}
H. Matsumura, \emph{Commutative ring theory}, Cambridge Studies in Advanced Math, \textbf{8}, (1986).

\bibitem{Mig}
J. Migliore, F. Zanello, \emph{The strength of the weak Lefschetz property}, Illinois J. Math.\textbf{ 52} (2008), no. 4, 1417--1433.

\bibitem{sal}J.D.
Sally,
\emph{Numbers of generators of ideals in local rings}, Marcel Dekker, Inc., New York-Basel, 1978.

\bibitem{sal2}J.D.
Sally,
\emph{Stretched Gorenstein rings}, J. London Math. Soc. \textbf{20} (1979), no. 2, 19–-26.

\bibitem{sch}
H. Schoutens, \emph{Absolute bounds on the number of generators of Cohen-Macaulay ideals of height two},
Bull. Belg. Math. Soc. Simon Stevin
Volume \textbf{13}, Number 4 (2006), 719--732.

\bibitem{trung}
N.V. Trung, \emph{Bounds for the minimum number of generators of generalized Cohen-
Macaulay ideals}, J. Algebra \textbf{90} (1984), 1-–9.

\bibitem{bus}J.
St\"{u}ckrad, and W. Vogel,  \emph{Buchsbaum rings and applications. An interaction between algebra, geometry, and topology}, Mathematische Monographien {\bf{21}}, VEB Deutscher Verlag der Wissenschaften, Berlin, 1986.

\bibitem{wat}
J. Watanabe, \emph{The Dilworth number of Artinian local rings and finite posets with
rank function}, Proceedings of Conference on Commutative Algebra and Combinatorics,
Kyoto 1985, Kinokuniya Co. and North-Holland Publ. Co.

\bibitem{wat2}
J. Watanabe, \emph{The Dilworth number of Artin Gorenstein rings}, Adv. Math. \textbf{76} (1989),  194–199.


\bibitem{bcan}J.S.
Okon, J.P. Vicknair,  \emph{A Gorenstein ring with larger Dilworth number than Sperner number}, Canad. Math. Bull. \textbf{43} (2000),  100-–104.

\bibitem{vas}
W. Vasconcelos, \emph{Cohomological degrees of graded modules}, Six lectures on Commutative Algebra (J. Elias, J.M. Giral, R.M. Miró-Roig, and S. Zarzuela, eds.), Progress in Math., vol. 166, Birkhäuser, 1998,  345–-390.
\end{thebibliography}

\documentclass[10pt,a4paper]{amsart}
\usepackage{amscd,amssymb,amsopn,amsmath,amsthm,graphics,amsfonts,enumerate,verbatim,calc}
\usepackage[dvips]{graphicx}

\usepackage{mathpazo}
\usepackage{color}
\usepackage{latexsym}
\usepackage{amsthm,amsfonts,amssymb,mathrsfs}
\usepackage{rotating}
\usepackage[leqno]{amsmath}
\usepackage{xspace}
\usepackage[all]{xy}
\usepackage{longtable}
\headsep=1cm \numberwithin{equation}{section}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{notation}[theorem]{Notation}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{fact}[theorem]{Fact}
\newtheorem{example}[theorem]{Example}
\newtheorem{observation}[theorem]{Observation}
\newtheorem{discussion}[theorem]{Discussion}
\newtheorem{question}[theorem]{Question}

\newtheorem{acknowledgement}{Acknowledgement}

\newcommand{\rr}{\operatorname{r}}
\newcommand{\D}{\operatorname{d}}
\newcommand{\hdeg}{\operatorname{hdeg}}
\newcommand{\Ass}{\operatorname{Ass}}

\newcommand{\soc}{\operatorname{soc}}

\newcommand{\rad}{\operatorname{rad}}

\newcommand{\pd}{\operatorname{p.dim}}

\newcommand{\Ext}{\operatorname{Ext}}

\newcommand{\Hom}{\operatorname{Hom}}
\newcommand{\gr}{\operatorname{gr}}

\newcommand{\depth}{\operatorname{depth}}

\newcommand{\emb}{\operatorname{emb dim}}
\newcommand{\e}{\operatorname{e}}
\newcommand{\Char}{\operatorname{char}}

\newcommand{\coker}{\operatorname{coker}}
\newcommand{\HH}{\operatorname{H}}

\newcommand{\m}{\operatorname{m}}

\newcommand{\lo}{\longrightarrow}
\newcommand{\fm}{\frak{m}}
\newcommand{\fp}{\frak{p}}
\newcommand{\fq}{\frak{q}}
\newcommand{\fa}{\frak{a}}

\newcommand{\fn}{\frak{n}}

\begin{document}

\author[]{mohsen asgharzadeh}

\address{}
\email{mohsenasgharzadeh@gmail.com}

\title[ ]
{on the Rees,     Dilworth and Matsumura number}

\subjclass[2010]{ Primary 13H10; 13E15}
\keywords{homologically exact; Rees  and Dilworth number; complete-intersection; Gorenstein rings}
\dedicatory{}
\begin{abstract}
We deal with some questions posted by Matsumura and Watanabe about the Rees and Dilworth number, and their higher-dimensional versions.
\end{abstract}

\maketitle

\section{Introduction}

Rings with uniform bound on the minimal number of generators of their ideals are so special. Sally proved that they are rings of Krull dimension
at most one (see \cite{sal}). Here, $\ell(-)$ is the length function and $\mu(-)$ stands for the minimal number of generators. Let $(R,\fm)$ be of Krull dimension
at most one. Then $\mu(\fa)\leq \ell(R/ x R)$ for any ideal $\fa$ and for any element $x\in \fm$. In particular,  $\rr(R):=\inf\{\ell(R/ x R): x\in\fm\}$ is an upper bound for $\D(R):=\sup\{\mu(\fa): \fa\subset\fm\}$. These numbers were introduced by J. Watanabe in the zero dimensional case and by Matsumura in the 1-dimensional case. They called them the Rees and the Dilworth number, respectively.
In particular, $\D(R)\leq \rr(R)$ (also see \cite{trung}). The  local  ring $(R,\fm)$ is called \textit{exact} if $\D(R)=\rr(R)$. This word was coined by Matsumura, but after that no one has used this word.
This sometimes is called the maximum-minimum property, see e.g. \cite{i}. In \cite[Question A]{M},  Matsumura asked the following question:

\begin{question}
Are zero-dimensional complete-intersections  exact?
\end{question}

By using a high-school computation (and also by a modern argument) we present a  non-exact local complete-intersection ring  over $\mathbb{F}_2$.  Section 3 deals with the following question of Watanabe:

\begin{question}\label{1.4}
	Can one construct an example of an artinian Gorenstein local ring, which is not
	exact over a field of characteristic zero?
\end{question}

By $\m(R)$  we  mean
$\inf\{\ell(R/\underline{x}R):\underline{x}\textit{ is a  parameter sequence}\}$ and we call it  the \textit{Matsumura number}. In \cite[Question C]{M},  Matsumura asked:

 \begin{question} \label{mat2} i) Is $\m(R)$ an interesting integer?\\
ii) Is there any higher dimensional generalization of the exactness?
\end{question}

In \S 4  we present some basic properties of $\m(R)$. Then we apply an idea taken from  Gunston's thesis,
to introduce a new family of noetherian local rings which we  call them \textit{homologically exact} with the property that
they are the same as of the exact rings provided they are 1-dimensional. The class  of homologically exact rings contains the class of Buchsbaum rings.
This drops a dimension-restriction from a result of Ikeda.

In the sequel all rings are commutative, noetherian and local. The symbol m stands for the maximal
ideal. The modules are finitely generated, otherwise specializes. The Matsumura's beautiful book \cite{Mat}
is our  reference book.
\section{A complete-intersection ring which is not exact }

In this section $R$ is the ring $\mathbb{F}_2[X,Y,Z]/(X^2,Y^2,Z^2)$.
 The ring $R=\bigoplus_{i=0}^3R_i$ is both graded and local.
In particular, any element $a$ of $R $ is of the form $a=a_0+a_1+a_2+a_3$ where $a_i\in R_i$.

\begin{lemma}\label{12}
Let  $a\in R_2+R_3$ and $0\neq b\in R_1$. Then $(0:_Ra+b)\nsubseteqq \fm^2$.
\end{lemma}

\begin{proof}
 We bring the following claim:
\begin{enumerate}
\item[Claim] A) Let  $r\in R$ be such that $r_0=0$. Then $r^2=0$. Indeed, $(R_2)^2=  (R_3)^2=0$.
It remains to assume that $r$ is homogeneous and  is of degree one. The claim is clear from  the above table.
\end{enumerate}
 Suppose on the contradiction that $(0:_Ra+b)\subseteq \fm^2$. From Claim A), one has $(a+b)^2=0$. Thus, $a+b\in(0:_Ra+b)\subseteq \fm^2$. Since $a\in R_2$  we have $b\in \fm^2$, a contradiction.
\end{proof}

The following easy (but important) fact plays an essential role in this paper:

\begin{fact}\label{main}(Watanabe, Matsumura,  Ikeda; see \cite[Theorem 2]{M})
Let $(A, \fm)$ be a
local Noetherian ring, let $M$ be a finitely generated A-module with $\dim M \leq 1$, let $N\subset M$ and  $a \in \fm$. Then  $\mu(N)\leq\ell(M/aM)$. Furthermore, $\mu(N)=\ell(M/aM)$ if and only if the following  claims hold:
\begin{enumerate}
\item[a)] $\ell(M/aM)<\infty$, and $\dim (M/N) = 0$.
\item[b)] $(0 :_M a) \subset N$, and
$\fm N = aN$.
\end{enumerate}
\end{fact}

\begin{example}
One has $\rr(R)=4$.
\end{example}

\begin{proof}
	  In the light  of \cite[Example 1.4]{i} we observe that $\D(R)=3$\footnote{One may prove this by hand as  a high school exercise (this is tedious at least for me).}.
Recall that  $\D(R)\leq \rr(R)$.  Hence, $\rr(R)\geq 3$. Since $\ell(R/xR)=4$, it is enough to show that $\inf\{\ell(R/ a R): a\in\fm\} \neq 3.$
Recall that $\mu(\fm^2)=\D(R)$. We apply Fact \ref{main} for $N:=\fm^2$ and $M:=R$. We are going to show that $(0 : r) \nsubseteqq \fm^2$ for all $r\in\fm$. We may assume that
$r\neq 0$ and that $r_0=0$.  Suppose first that $r$ is homogeneous. To this end, we bring 3 claims according to $\deg(a)$.
 \begin{enumerate}
\item[Claim]A)
Let  $r\in R_1$. Then $(0:_Rr)\nsubseteq \fm^2$. \\Indeed,
due to Claim A) from Lemma \ref{12}, $r^2=0$. So, $r\in (0:_Rr)\setminus \fm^2$.
\item[Claim]B)
Let  $r\in R_2$. Then $(0:_Rr)\nsubseteq \fm^2$.
Indeed,
suppose $r:=xy+yz+zx$. Then $$(x+y)r=x^2y+xyz+zx^2+xy^2+y^2z+xzy=2xyz=0.$$ So, $(x+y)\in (0:_Rr)\nsubseteq \fm^2$.
Suppose $r:=xy+yz$. Then $yr=0$. So, $y\in (0:_Rr)\nsubseteq \fm^2$. By the symmetric, the same claim hold for both of  $xy+xz$, and $yz+xz$.
Suppose $r:=xy$. Then $yr=0$. So, $y\in (0:_Rr)\nsubseteq \fm^2$. By the symmetric, the same claim hold for both of  $xy$, and $xz$.
Thus, the desired claim checked for all elements of $R_2$.
\item[Claim]C)
Let  $r\in R_3$. Then $(0:_Rr)\nsubseteq \fm^2$.\\
Indeed,
we have $R_i=0$ for all $i>3$. Thus $R_1r=0$. So, $R_1\subset (0:_Rr)\nsubseteq \fm^2$.
\end{enumerate}

Thus, the proof in the homogenous case follows.
Write $r=r_1+r_2+r_3$. In view of Lemma \ref{12}, we can assume that $r_1=0$, i.e., $r=r_2+r_3$. Since $r$ is not homogeneous, $r_2$ and $r_3$
are nonzero. By the proof of Claim B) there is non-zero $b\in R_1$ such that $r_2b=0$.
Recall that $R_1R_3=0$, because $R_i=0$ for all $i>3$.
In particular, $bR_3=0$. So, $b\in(0 : r_2+r_3)$. Combine this along with  $b\notin \fm^2$ to get that $(0 : r_2+r_3) \nsubseteqq \fm^2$.
\end{proof}

\begin{corollary}\label{co}
The ring $R$ is a zero-dimensional local  complete-intersection ring which is not exact.
\end{corollary}

One can prove $\rr(R)\geq4$ by using the machinery of \textit{Lefschetz} properties. Let us recall the following  comment  of Watanabe:

\begin{discussion} \label{dis}
Let $(A,\fm)$ be a graded artinian algebra over a field $K$.
\begin{enumerate}
\item[i)] Recall that $A$ has the weak Lefschetz property if
there is an element  $\ell$ of degree $1$ such that the multiplication $×\ell:A_i \to A_{i+1}$ has maximal
rank, for every $i$.
\item[ii)]The Hilbert function
$(h_0,\ldots,h_c)$ is said to be unimodal if the  there exists $j$ such that
$$h_0\leq h_1\leq h_i \leq \ldots\leq h_j \geq h_{j+1}\geq\ldots \geq h_{c-1}\geq h_{c}$$
\item[iii)](See \cite[Proposition 3.5]{wat3}) Suppose $A$  has
unimodal Hilbert function, and let $\ell \in A_1$. Then $$\ell\textit{  is a weak Lefschetz element for }A\Longleftrightarrow\sup\{\dim A_i\}=\ell(A/\ell A).$$
\item[iv)] Due to \cite[Lemma 2.2]{i}, we have $\rr(A)=\inf\{\ell(A/r):r\in A_1\}.$
\item[v)] (See \cite[Theorem 13]{iw}) Set $A:=\frac{K[X_1, \ldots , X_n]}{(X_1^{d_1},\ldots,
X_n^{d_n})}$. Then the maximum and the minimum
of $\{I \lhd A | \mu(I) = \D(A)\}$ are powers of the maximal ideal. In particular, $\sup\{\mu(\fm^i)\}= \D(A)$.
\item[vi)] It may be $\sup\{\mu(\fm^i)\}\neq \D(A)$ (see \cite[Example 4.4]{i}).
\end{enumerate}
\end{discussion}

Now, we are ready to present a  modern proof of Corollary \ref{co}:

\textbf{Second proof of Corollary \ref{co}}.
One has $\rr(R)\geq4$: Indeed, one can check easily that $R$ has no Lefschetz element (see e.g. \cite[Remark 2.9]{wja}). The Hilbert function of $R$
is unimodal. In view of Discussion \ref{dis}(iii) $\sup\{\dim R_i\}\neq\inf\{\ell(R/\ell R):\ell\in R_1\}$. By Discussion \ref{dis}(iv) $\rr(R)=\inf\{\ell(R/rR):r\in R_1\}$. In the light of Discussion \ref{dis}(v) we see $\D(R)=\sup\{\dim R_i\}$. Also, $\D(R)\leq \rr(R)$. Putt all of these together to observe
$\rr(R)>\sup\{\dim R_i\}=3.$

\section{Dealing with Question \ref{1.4}}

\begin{lemma}\label{pro}
Let $R$ be  an artinian local ring such that $\ell(R)\leq\mu(\fm)+2$. Then $R$ is exact.	
	\end{lemma}

\begin{proof} The case $\ell(R)=\mu(\fm)$ never happens, because $\mu(\fm)=\ell(\fm/\fm^2)<\ell(R)$. Suppose first that $\ell(R)=\mu(\fm)+1$. Let $x\in\fm$ be nonzero.
Then $$\mu(\fm)\leq\sup\{\mu(\fa)\}=\D(R)\leq\rr(R)=\inf\{\ell(R/rR)\}\leq\ell(R/xR)\leq\ell(R)-1=\mu(\fm).$$
Therefore, $\D(R)=\rr(R)=\mu(\fm)$. So, $R$ is exact.

Now, we deal with $\ell(R)=\mu(\fm)+2$. Since $\mu(\fm)=\ell(\fm/\fm^2)$
	we may assume that $ \fm^2\neq 0$.  Let $x\in\fm\setminus \fm^2$. There is an $y\in\fm$ such that $xy\neq 0$. Indeed, if such a thing does not happen, then it yields that $\fm^2=0$ which is excluded.
	Clearly, $\ell(R/xR)<\ell(R)=\mu(\fm)+2.$ We claim that $\ell(R/xR)=\mu(\fm)$.
	Indeed, recall   that $\ell(R/xR)\geq\rr(R)\geq \D(R)\geq\mu(\fm)$. If $\ell(R/xR)\neq\mu(\fm)$, then $\ell(R/xR)=\mu(\fm)+1$. Also,  $\ell(R/xyR)\leq\ell(R)-1$. Since $R/xyR\to R/xR\to 0$, $$\mu(\fm)+1=\ell(R)-1\geq\ell(R/xyR)\geq\ell(R/xR)=\mu(\fm)+1.$$ So, $\ell(R/xyR)=\ell(R/xR)$. Now, we look at the exact sequence $$0\lo xR/yxR\lo R/xyR\lo R/xR\lo 0.$$Taking length, implies that
	$xyR=xR$. By Nakayama's lemma, $x=0$. This contradiction says that $\ell(R/xR)=\mu(\fm)$. By definition, $R$ is  exact.
	\end{proof}

\begin{lemma}\label{m2}
	Let $R$ be an artinian local ring such that $\fm^2=0$. Then $R$ is exact.	
\end{lemma}

\begin{proof} We may assume that $\fm$ is nonzero. First note that $\mu(\fm)=\ell(\fm/\fm^2)=\ell(\fm)$.
	In view of  $0\to \fm \to R\to R/\fm\to 0$ we get that
	$\mu(\fm)=\ell(\fm)=\ell(R)-1$.  The claim is now clear from
	Lemma \ref{pro}.
\end{proof}

\begin{proposition}\label{3}
	Let $R$ be an artinian Gorenstein local ring such that $\fm^3=0$. Then $R$ is exact.	
\end{proposition}

\begin{proof} By Lemma \ref{m2}, we may assume that $\fm^2$ is nonzero. Note that $\mu(\fm^2)=\ell(\fm^2/\fm^3)=\ell(\fm^2)$. Since
	$R$ is Gorenstein  its socle is one-dimensional.
Due to $\fm^3=0$, we have $\fm^2\subset \soc(R)$. Consequently, $1\leq\ell(\fm^2)\leq\ell(\soc(R))=1$.
 Thus,
	$\mu(\fm^2)=\ell(\fm^2)=1$. In view of  $0\to \fm^2 \to \fm\to \fm/\fm^2\to 0$ we get that
	$\mu(\fm)=\ell(\fm)-\ell(\fm^2)=\ell(\fm)-1$.  Also, we apply the exact sequence  $0\to \fm  \to R\to R/\fm \to 0$ to see
  $\ell(\fm)=\ell(R)-1.$  Therefore, $\mu(\fm)=\ell(\fm)-1=\ell(R)-2.$ It remains to apply
	Lemma \ref{pro}.
\end{proof}
	
The Gorenstein assumption  in the above proposition  is important:
\begin{example}\label{e9m}
Let $R:=\frac{\mathbb{C}[X,Y,Z,W]}{ (X^2,Y^2,Z^2,W^2,XY,ZW)}$. In view of \cite[Page 160]{M} $R$ is not exact. Let us
present a proof: By a result of Watanabe if a monomial algebra $k[x_i]/I$ were be exact then we should have $(\sum x_i)\fm=\fm^2$ (this criterion works only for monomial algebras,  see Example \ref{ew}).
One has $\fm^2=(wx,wy,xz,yz)$ and 
$\fm(x+y+z+w)=(xz+xw,yz+yw,xz+yz,xw+yw)$. From this we see that $$\dim_{\mathbb{Q}}(xz+xw,yz+yw,xz+yz,xw+yw)_2<4=\dim_{\mathbb{Q}}(\fm^2)_2.$$We apply Watanabe's criterion to see $R$ is not exact.
Note that  $\fm^3$ is generated by degree three monomials in $\{x,y,z,w\}$. They are zero by the relations $x^2=y^2=z^2=w^2=xy=zw=0$. So,
$\fm^3=0$.\end{example}

\begin{notation} \label{not}Let $(R,\fm,k)$ be a local  artinian $k$-algebra. There is $\ell$ such that $\fm^{\ell+1}=0$ and $\fm^{\ell}\neq0$.
 We set $h_i:=\dim_k(\frac{\fm^i}{\fm^{i+1}})$. By the Hilbert function of $R$
we mean $\HH(R):=(1,h_1,\ldots,h_{\ell})$.
\end{notation}

\begin{fact}\label{low}(Watanabe)
	Let $(R,\fm,k)$ be an artinian  local ring. The following assertions hold.
	\begin{enumerate}
		\item[i)] If $\emb(R)=1$, then $R$ is exact.
		\item[ii)] If $\emb(R)=2$ and $k$ is infinite, then $\D(R)=\rr(R)$.
\item[iii)] If $\rr(R)\leq3$ and $k$ is infinite, then $\D(R)=\rr(R)$.
\item[iv)] One has $\rr(R)\leq \rr(\gr_{\fm}(R))$\end{enumerate}
\end{fact}

\begin{proof}	i) Let $x$ be such that $\fm=(x)$. Then $\mu(\fm)=\ell (R/xR)$. Thus, $\D(R)=\rr(R)=1$.

ii) This is in \cite[Theorem 4.2]{i}.

iii) If $\D(R)\leq2$, by ii) we get the claim. We may assume that $\D(R)=3$. Recall that $\D(R)\leq\rr(R)$. So, $\D(R)=\rr(R)=3$.

iv)  This is in \cite[Proposition 2]{wat2}.
\end{proof}

Following Sally \cite{sal2}, a local artinian ring  is said to be \textit{stretched} if $\fm^2$ is a principal ideal.

\begin{proposition} \label{st}
Any  stretched Gorenstein algebra is exact.
\end{proposition}

\begin{proof}
We are going to use the associated graded ring   $\gr_{\fm}(R)$. This is a graded ring and is of zero dimension (see \cite[Theorem 13.9]{Mat}). Keep the notation be as of  Notation \ref{not}. By Macaulay's theorem  $h_i=1$ for all  $2\leq i\leq \ell$
(Sally  \cite[Theorem 1.1(ii)]{sal2} proved this without any use of Macaulay's theorem).
Thus,  $\HH(R)=(1,n,1,\ldots,1)$. Any nonzero map
from (resp. to) a 1-dimensional vector space is injective (resp. surjective). Conclude by this that  $\gr_{\fm}(R)$ has a weak Lefschtez element.
In the light of Discussion  \ref{dis}, $\rr(\gr_{\fm}(R))=n$. By  Fact \ref{low}(iv), $\rr(R)\leq \rr(\gr_{\fm}(R))$. Therefore,
$$n=\mu(\fm)\leq\D(R)\leq\rr(R)\leq\rr(\gr_{\fm}(R))=n.$$So, $R$ is exact.
\end{proof}

\begin{proposition}\label{als}
Any almost stretched Gorenstein algebra $(A,\fm)$ over an algebraically closed field of characteristic $0$ is exact.
\end{proposition}

\begin{proof}
Due to Fact \ref{low} we  may assume that $n:=\emb(A)\geq3$.
Let us clarify the assumption: There is  a regular local ring $(R,\fn)$ such that $A = R/I$. By the almost stretched
we mean that $\mu(\fm^2)=2$. Then $\HH_A(i)=2$  for $2\leq i\leq t$ and $\HH_A(i)=1$ for all $t+1\leq i\leq s$.
We recall the following classification result of Elias and Valla (see \cite[Theorem 4.1]{ev}):
There is a  basis
$Y_1,\ldots,Y_n$ of $\fn$ and $b\in R$ such that $I$ is generated by the following four classes of functions:

\begin{enumerate}
\item[i)] $Y_iY_j$ with $1\leq i < j \leq n$ and $(i, j)\neq (1, 2)$,
\item[ii)] $Y^2_j - d_jY^s_1$ with $3\leq j \leq n$,
\item[iii)] $Y^2_2 - b Y_1 Y_2 -cY^{s-t+1}_1$,
\item[iv)] $Y^t_1Y^2$,
\end{enumerate}where $\{d_j,c\}$ are invertible.
Note that $\fm^2=(y_iy_j)_{i,j}$. From the first three relations we see that $\fm^2=(y_1^2,y_1y_2)$. Let $\xi:=y_1+y_3$. Then $y_1\xi=y_1^2$ and $y_2\xi=y_1y_2$.
Thus, $\fm\xi=\fm^2$. Clearly, $(0:\xi)\subset\fm$.
In the light of Fact  \ref{main}, $\ell(A/ \xi A)=\mu(\fm)=n$. Therefore, $n\leq\D(A)\leq\rr(A)\leq n$, which is what we want to prove.
\end{proof}

\begin{corollary}\label{77}
Let $(R,\fm,k)$ be a Gorenstein artinian local  $k$-algebra, where $k=\overline{k}$ is of zero-characteristic. If $\ell(R)\leq 7$, then $R$ is exact.
\end{corollary}

\begin{proof}
First,  assume that $\ell(R)= 6$. In this case $\mu(\fm)\leq3$ (see Lemma \ref{pro}). The case $\mu(\fm)\leq2$ is in Fact \ref{low}.
Then $\mu(\fm)=\ell(\fm/ \fm^2)=3$. In view of  $0\to \fm^2 \to \fm\to \fm/\fm^2\to 0$, we observe $$\ell(\fm^2)=\ell(\fm)-3=5-3=2.$$ Either $\mu(\fm^2)=1$ or $\mu(\fm^2)=2$.  Since $R$ is Gorenstein,
$(1,3,2)$ is not admissible. Conclude by this that $\HH(R)=(1,3,1,1)$, i.e., $R$ is stretched. So, $R$ is exact.
Without loss of the generality we can assume that $\ell(R)= 7$. In view of Proposition \ref{3} we may and do assume that $\fm^3\neq 0$. Also, in the light  of Lemma \ref{pro} we may assume that
$\mu(\fm)\leq 4$. There are the following two possibilities:

\begin{enumerate}
\item[i)] One has $\mu(\fm)=4$ (in this case the claim holds without Gorenstein assumption). 	
Indeed, there are $x,y$ and $z$ in $\fm$ such that $xyz\neq0$. Note that
$\ell(R/(xyz))\leq6$. By Nakayama, $xyR/xyzR\neq 0$. Since $0\to xyR/xyzR\to R/xyzR\to R/xyR\to 0$, we have $\ell(R/(xy))<\ell(R/(xyz))\leq6$.
Similarly, $$\ell(R/(x))<\ell(R/(xy))<\ell(R/(xyz))\leq6.$$ Hence, $\ell(R/(x))\leq 4$. Recall that $4\leq\mu(\fm)\leq\ell(R/(x))\leq 4$, i.e., $R$ is exact.

\item[ii)] One has $\mu(\fm)\leq3$. Indeed, by Fact \ref{low}  $\mu(\fm)=3$.
Suppose that $\HH(R):=(1,3,1,1,1)$.  Let us check this case directly. Let $x,y,z,w$ be such that $xyzw\neq 0$ (note that $\fm^4\neq 0$). We deduce from $$\ell(R/(x))<\ell(R/(xy))<\ell(R/(xyz))<\ell(R/(xyzw))\leq6$$
that $\ell(R/(x))\leq 3$. Recall that $3\leq\mu(\fm)\leq\ell(R/(x))\leq 3$. Therefore, $R$ is exact.
By Macaulay's theorem, $(1,3,1,2)$ is not admissible.
The only remaining case is $\HH(R)=(1,3,2,1)$. By definition, $R$ is almost stretched. So, $R$ is exact.
\end{enumerate}
\end{proof}

\begin{remark}
Recall that $(1,3,1,2)$ is not admissible.  Here, we derive it as an application of the theory of exact rings. Suppose it is an $O$-sequence of  a ring $R$. Let $A:=\gr_{\fm}(R)$ and let $x\in A_1$ be such that $x A_2\neq 0$.
We can pick this, because $A_3\neq0$. Note that
$xA=xA_0\bigoplus x A_1\bigoplus  xA_2$.  Clearly, $\dim_k(xA_0)=1$. We see $0\neq xA_1\subset A_2$. Combine this along with  $\dim_k(A_2)=1$
to observe that $\dim_k(xA_1)=1$. Since $0\neq xA_2\subset A_3$ and that $\dim_k(A_2)=1$ we get that $\dim_k(xA_2)=1$.
We apply these to conclude that
$\dim_k(xA)=\sum_{i=0}^2\dim_k(xA_i)=1+1+1=3.$  We look at the exact sequence $0\to xA\to A\to A/xA\to 0$ to deduce
$\ell(A/xA)=\ell(A)-\ell(xA)=7-3=4.$ This proof shows that $x$ is a weak Lefschetz element. By Discussion  \ref{dis} $\ell(A/xA)=\sup\{\mu(A_i)\}=3$. This contradiction says that
$(1,3,1,2)$ is not admissible.
\end{remark}

The  ring presented in \S 2 was of length $8$. Despite of this we have:

\begin{corollary}\label{8}
Let $(R,\fm,k)$ be a Gorenstein artinian local  $k$-algebra, where $k=\overline{k}$ is of zero-characteristic. If $\ell(R)= 8$, then $R$ is exact.
\end{corollary}

\begin{proof} 
By Fact \ref{low} we may assume $\mu(\fm)>2$. In view of Proposition \ref{3} we can assume that $\fm^3\neq 0$. Also, in the light  of Lemma \ref{pro} we may assume that
$\mu(\fm)\leq 5$. Since $R$ is Gorenstein, the final coordinate of its  Hilbert function is one. The following are the possible shape of Hilbert functions:
$$\{(1,6,1),(1,5,1,1),(1,4,2,1),(1,4,1,1,1),(1,3,2,1,1),(1,3,3,1),(1,2,\ldots,1)\}.$$
 The  stretched and  almost stretched cases were handled in Proposition \ref{st} and \ref{als}. Then we may assume that $\HH(R)=(1,3,3,1)$.
We are going to apply the associated graded ring   $\gr_{\fm}(R)$. This is a graded ring with Hilbert function $(1,3,3,1)$.
 Watanabe   proved that if the Hilbert function of a  local ring  is symmetric then the associated graded ring $\gr_{\fm}(-)$ is  Gorenstein (see \cite[Proposition
9]{wat2}).   The assumptions
	imply that $\gr_{\fm}(R)\simeq k[x,y,z]/ I$ where $I$  is Gorenstein, graded and of height three.
	Since  $h_3(k[x,y,z]/ I)=1$, $\gr_{\fm}(R)$ has weak Lefschetz property (see \cite[Corollary 3.7]{Mig}). Combine this along with Discussion \ref{dis}(iii) to observe that $$\rr(\gr_{\fm}(R))=D(\gr_{\fm}(R))=\sup\{h_i(\gr_{\fm}(R))\}=3.$$ We apply   Fact \ref{low}(iv) to deduce $\rr(R)\leq \rr(\gr_{\fm}(R))=3$. Recall that $$3=\mu(\fm)\leq\D(R)\leq\rr(R)\leq 3.$$  So, $\D(R)=\rr(R)$.
The proof is now complete.
\end{proof}

The non-exact ring presented in Example \ref{e9m} is of length 9 and it contains $\mathbb{C}$ (it is not  Gorenstein). Despite of this, there is a complete list of artinian Gorenstein $k$-algebra of length $9$. The possible
Hilbert functions are $$\{  (1,2,\ldots,1),(1,3,2,2,1),(1,3,3,1,1),(1,4,3,1),(1,4,2,1,1),(1,5,2,1),
(1,5,1,1,1),(1,6,1,1),(1,7,1)\}.$$The cases $\fm^2=0$, stretched, and almost stretched are exact. Thus, the only nontrivial
cases are $(1,3,3,1,1)$ and $(1,4,3,1)$.

\begin{fact}\label{93}(See \cite[\S 2]{ca})
Let $(R,\fm,k)$ be a Gorenstein artinian local  $k$-algebra, where $k$ is of zero-characteristic. If $\HH(R)=(1,3,3,1,1)$, then $R:=\frac{k[x_,x_2,x_3]}{I}$ where $I$ is of the form:
\begin{enumerate}
\item[i)]  $(x_1x_2 +x^2_3,x_1x_3,x_2^2 -tx^2_3 -x^{\ell}_1)$
\item[ii)] $ (x_1x_2,x_1x_3,x_2x_3,x^3_2 -x_1^{\ell},x^3_3 -x_1^{\ell})$
\item[iii)] $(x_1x_2,x_2x_3,x^2_1,x_1x^2_3 -x_2^{\ell},x^3_3-x_2^{\ell})$
\item[iv)] $(x_1x_2,x_2x_3,x^2_1 -x^{\ell-1}_3,x_1x^2_3,x^3_2 -x_3^{\ell})$
\item[v)] $(x_1x_2 -x^{\ell-1}_3 ,2x_1x_3 +x^2_2,x^2_1,x_1x^2_3,x_2x^2_3)$
\item[vi)] $(x_1x_2,x^2_1 -x^{\ell-1}_3,x^2_2 -x^{\ell-1}_3,x_1x^2_3,x_2x^2_3 )$.
\end{enumerate}
\end{fact}

\begin{example}
Let $R:=\frac{k[x_,x_2,x_3]}{(x_1x_2,x_2x_3,x^2_1,x_1x^2_3 -x_2^{\ell},x^3_3-x_2^{\ell})}$. Then $R$ is exact.
\end{example}

\begin{proof}
It is easy to see that $\fm^2=(x_2^2,x_3^2,x_1x_3)$. Set $\xi:=x_1+x_2+x_3$. Then
$$\fm\xi=(x_1\xi,x_2\xi,x_3\xi)=(x_1x_3,x_2^2,x_3x_1+x_3^2)=(x_2^2,x_3^2,x_1x_3)=\fm^2.$$
In view of Fact \ref{main}, $R$ is exact.
\end{proof}

\begin{fact}\label{94}(See \cite[\S 5]{ca})
Let $(R,\fm,k)$ be a Gorenstein artinian local  $k$-algebra, where $k$ is of zero-characteristic. If $\HH(R)=(1,4,3,1)$, then $R:=\frac{k[x_,x_2,x_3,x_4]}{I}$ where $I$ is of the form:
\begin{enumerate}
\item[i)]  $ \left(x_1x_2 +x^2_3,x_1x_3,x^2_2 -ax^2_3-x_1^{\ell-1},x_ix_4,x_4^2-x_1^{3}\right)_{i<4}$ for some $a\in k$,
\item[ii)] $\left(x^2_1,x^2_2,x^2_3 +2x_1x_2,x_ix_4,x_4^2 -x_1x_2x_3\right)_{i<4}$
\item[iii)] $\left(x^2_1,x^2_2,x^2_3,x_ix_4,x^2_4 -x_1x_2x_3\right)_{i<4}$
\item[iv)] $\left(x_1x_2,x_1x_3,x_2x_3,x^3_2 -x^3_1,x^3_3 -x^3_1,x_ix_4,x^2_4-x^3_1\right)_{i<4}$
\item[v)] $\left(x^2_1,x_1x_2,x_2x_3,x^3_2 -x^3_3,x_1x^2_3 -x^3_3,x_ix_4,x^2_4 -x^3_3\right)_{i<4}$
\item[vi)] $\left(x^2_1,x_1x_2,2x_1x_3 +x^2_2,x^3_3,x_2x^2_3,x_ix_4,x^2_4 -x_1x^2_3\right)_{i<4}$.
\end{enumerate}
\end{fact}

\begin{example}
Let $R:=\frac{k[x_,x_2,x_3,x_4]}{(x^2_1,x^2_2,x^2_3,x_ix_4,x^2_4 -x_1x_2x_3)}$. Then $R$ is exact provided $\Char (k)\neq 2$.
\end{example}

\begin{proof}
It is easy to see that $\fm^2=(x_1x_2,x_1x_3,x_2x_3)$.
Set $\xi:=x_1+x_2+x_3$. Then $\fm\xi:=(x_1x_2+x_1x_3,x_1x_2+x_2x_3,x_1x_3+x_2x_3)$. Since $\Char (k)\neq 2,$ 
$$x_1x_2=\frac{1}{2}(x_1x_2+x_1x_3)+\frac{1}{2}(x_1x_2+x_2x_3)-\frac{1}{2}(x_1x_3+x_2x_3).$$
Thus, $x_1x_2\in\fm\xi$. Therefore, $\fm\xi=(x_1x_2,x_1x_3,x_2x_3)=\fm^2$.\footnote{If $\Char (k)=2$, then $\fm\xi\subsetneqq\fm^2$. Indeed, $x_1x_2+x_1x_3=(x_1x_2+x_2x_3)+(x_1x_3+x_2x_3)$. Therefore,
$\dim_{k}(\fm\xi)_2<3=\dim_{k}(\fm^2)_2$. From this $\fm\xi\subsetneqq\fm^2$.} In view of Fact \ref{main}, $R$ is exact.
\end{proof}

Here, we study exactness of Gorenstein rings with $\fm^4=0$  in some examples:

\begin{example} \label{ew1}
Let $A:=\mathbb{Q}[u, v, x, y,z]$ and let $I:=\left\{f\in A|f(\partial_u,\partial_v,\partial_x,\partial_y,\partial_z\right)(u^2x + uvy + v^2z)=0\}.$
Then $R:=A/I$ is an  exact Gorenstein ring.
\end{example}

\begin{proof}By a result of Watanabe the Hilbert series
of $R$ is $1+5t+5t^2+1$. Note that $$(2\partial_u\partial_y-\partial_v\partial_z)(u^2x + uvy + v^2z)=0.$$ Also,
$(\partial_u\partial_x-2\partial_v\partial_y)(u^2x + uvy + v^2z)=0$. These mean that $2uy-vz\in I$ and $ux-2vy\in I$.
Let $$J:=\langle z^2, yz, xz ,uz, y^2, xy ,2uy-vz, x^2, vx, ux-2vy, v^3, uv^2, u^2v, u^3\rangle.$$
It is easy to see that $J\subset I$ and that the Hilbert  series
of $A/J$ is $1+5t+5t^2+1$. We look at the exact sequence of graded modules $0\to I/J\to A/J\to R\to 0$. By using the additivity of Hilbert function,
$I/J=0$. In particular, $J= I$ and that $R$ is local.
Let $\fa:=( x, y, z, uv, u^2, v^2)$. Then $(0:_Ru)\subset \fa$ and $\fm \fa=u\fa$. In the light of Fact  \ref{main}, $\mu(\fa)=\ell(R/uR)$.
Therefore, $\D(R)=\rr(R)=6$.
\end{proof}

\begin{example} \label{ew}
Let $A:=k[X_1, \ldots, X_5]$ be the polynomial ring over any field $k$ of zero-characteristic.
Let $I$ be the ideal generated by the
following quadric relations:
$$\Delta:=\{X_1X_3 + X_2X_3, X_1X_4 + X_2X_4, X_3^2 + X_1X_5 - X_2X_5,
X^2_4 + X_1X_5 - X_2X_5, X^2_1, X_2^2
, X_3X_4, X_3X_5, X_4X_5, X_5^2\}.$$
Let $R := A/I$. The ring $R$ is   Gorenstein and $\fm^4=0$. The following holds:\begin{enumerate}
\item[i)] One has $\ell(R/\sum_{i=1}^5 x_i)\neq\inf\{\ell(R/rR)\}$.
\item[ii)] The ring $R$ is exact. In fact, there are infinitely many
$r$ such that $\ell(R/rR)=\rr(R)=\D(R)=5$.
\end{enumerate}
\end{example}

\begin{proof}
In view of \cite{ga}, $R$ is a local Gorenstein ring with Hilbert series $\HH(t) = 1+5t+5t^
2+t^3$. Hence $\fm^4=0$.
Also, as a $k$-vector space, it has a basis consisting of $$\Gamma:=\{
1, x_1, x_2, x_3, x_4, x_5, x_1x_2, x_1x_3, x_1x_4, x_1x_5, x_2x_5, x_1x_2x_5 \}.$$The relations
$x_3^2 + x_1x_5 - x_2x_5=
x^2_4 + x_1x_5 - x_2x_5=0$ implies that $x^2_3=x_4^2$ and $x^2_4 + x_1x_5 =x_2x_5$.
Let $$\xi:=x_1+Ax_2+Bx_3+Cx_4+Dx_5\in R_1$$ be an element.
We are going to force conditions for which $\fm\xi=\fm^2$.
Let us compute $x_i\xi:$
\begin{enumerate}
\item[1)] $x_1\xi=Ax_1x_2+Bx_1x_3+Cx_1x_4+Dx_1x_5$,
\item[2)] $x_2\xi =x_1x_2+Bx_2x_3+Cx_2x_4+Dx_2x_5=x_1x_2-Bx_1x_3-Cx_1x_4+Dx_1x_5+Dx_3^2$,
\item[3)] $x_3\xi=x_3x_1+Ax_2x_3+Bx_3^2=(1-A)x_3x_1+Bx_3^2$,
\item[4)]  $x_4\xi=x_4x_1+Ax_2x_4+Cx_4^2=(1-A)x_1x_4+Cx_3^2$,	
\item[5)] $x_5\xi=x_5x_1+Ax_2x_5=(A+1)x_1x_5+Ax_3^2$.
\end{enumerate}
 We now  are ready to prove the desired claims:

\begin{enumerate}
\item[i)] Let $A=B=C=D=1$. We have $\fm\sum_{i=1}^5 x_i\neq \fm^2$, because $\dim_k(\fm\sum_{i=1}^5 x_i)_2<5=\dim_k(\fm^2)_2$.

\item[ii)]  Let $A:=-1$. We claim that $\fm\xi=\fm^2$ for any nonzero $B$, $C$ and $D$. Indeed,
from $5)$ we get that $x_3^2\in \fm\xi$. In view of $4)$ and  $3)$ we see $\{x_3x_1,x_1x_4\}\subset\fm\xi$.
Look at
\[\begin{array}{ll}
x_1\xi+x_2\xi
=(A+1)x_1x_2+(2D)x_1x_5+Dx_3^2=(2D)x_1x_5+Dx_3^2.
\end{array}\]Conclude by this that  $x_1x_5\in\fm\xi$.
In the light of  $x_3^2 + x_1x_5 - x_2x_5=0$ we see $x_2x_5\in\fm\xi$. Since
$\fm^2=\langle x_1x_2,x_3x_1,x_1x_4,x_2x_5,x_1x_5\rangle$ we have
$\fm\xi=\fm^2$. This says that $\xi$ is a Lefschetz element. The Hilbert function is unimodal. Combine these along with  Discussion \ref{dis}(iii) to conclude
that $\rr(R)=5$. Recall that $5\leq\D(R)\leq\rr(R)=5$. \end{enumerate}This is what we want to prove.
\end{proof}

\section{Higher dimensional exactness }

Suppose $R$ is  homomorphic image of a Gorenstein ring $A$ of same dimension  as of $R$ (e.g. $R$ is complete). Let $M$ be a finitely generated $R$-module of dimension $d$. If $d=0$ set $\D_h(M):=\D(M)$.
Note that $\dim(\Ext^{d-i}_A(M,A))\leq i$. Inductively,  define the \textit{homological Dilworth number} of $M$ by $\D_h(M):=\sum_{i=0}^{d-1}(^{d-1}_{i})\D_h(\Ext^{d-i}_A(M,A))+\e(M)$.
This is independent of the choose of  $A$ (see \cite[Proof of 3.9(a)]{vas}).

\begin{fact}\label{G}(Gunston \cite{dil})
Adopt the above notation. If $\dim R=1$, then $\D(R)=\D_h(R)$.
\end{fact}

\begin{lemma}\label{comp}
Adopt the above notation. Then
$\D_h(M) =\D_h(M\otimes_R\widehat{R})$.
\end{lemma}

\begin{proof}
	The proof is by induction on $d:=\dim M$. Since   finite-length module are complete, we may assume $d>0$. This is well-known that $\e(M)=\e(\hat{M})$. Recall that $R$ is  homomorphic image of a Gorenstein ring $(A,\fm_A)$ of same dimension  as of $R$.
Let $i<d$. By induction,  $\D_h(\Ext^{d-i}_A(M,A)\otimes_R\widehat{R})=\D_h(\Ext^{d-i}_A(M,A))$. Note that $\fm_A$-adic completion of a finitely generated $R$-module
$N$ is the same as of its $\fm_R$-adic completion,
when we view $N$ as an $A$-module via the map $A\to R$.  Then, $\Ext^{d-i}_A(M,A)\otimes_R\widehat{R}\simeq\Ext^{d-i}_{\hat{A}}(\hat{M},\hat{A})$.  Therefore, $\D_h(\Ext^{d-i}_A(M,A))=\D_h(\Ext^{d-i}_{\hat{A}}(\hat{M},\hat{A}))$. By definition, $\D_h(M)=\D_h(\hat{M})$.
\end{proof}
If $R$ is not  homomorphic image of a Gorenstein ring,  define
$\D_h(M) :=\D_h(M\otimes_R\widehat{R})$.  By the above lemma, this is well-define.
By a \textit{formal} module we mean a finitely generated $\widehat{R}$-module.
Also,  a formal module $\mathcal{M}$ is  called \textit{algebraic} if there is a finitely
generated $R$-module $M$ such that $\widehat{M}\simeq \mathcal{M}$.

\begin{fact}\label{algart}
i) Any finite length $\widehat{R}$-module  $\mathcal{M}$ is algebraic.

ii) Let $M$ be an $R$-module such that $\ell_R(\widehat{M})<\infty$. Then  $\ell_R(M)=\ell_{\widehat{R}}(\widehat{M})$.
\end{fact}

\begin{proof}
This follows by an easy induction and we left the routine details to the reader.
\end{proof}

\begin{corollary}\label{mc}
Let  $\mathcal{I}\vartriangleleft \widehat{R}$ be a full parameter ideal in a local ring $\widehat{R}$.
Then there is a full parameter ideal $I$ in $R$ such that $\ell(R/I)=\ell(\widehat{R}/\mathcal{I})$.
In particular, $\m(R)=\m(\widehat{R})$.
\end{corollary}

\begin{proof}
	By Fact \ref{algart}(i), there is a finitely generated $R$-module $M$ such that
$\widehat{M}\simeq\widehat{R}/\mathcal{I}$. We claim that $M$ is cyclic. Indeed, we look at the following exact sequence of formal modules $0\to \mathcal{I}\to \widehat{R}\stackrel{\mathcal{F}}\to \widehat{R}/\mathcal{I} \to 0.$
Since $\Hom_{\widehat{R}}(\widehat{R},\widehat{R}/\mathcal{I})\simeq\Hom_{R}(R,M)\otimes \widehat{R}$ there is an  $F\in \Hom_{R}(R,M)$
such that $\widehat{F}=\mathcal{F}$.
 Due to faithfully flatness of completion, $\coker(F)=0$, i.e., $F$ is surjective. So $M$ is cyclic. Put $I:=\ker F\lhd R$. There is a natural map $I\otimes \widehat{R}\to \mathcal{I}$.
We apply 5-lemma to observe that   $\widehat{I}\simeq\mathcal{I}$.
By Fact \ref{algart}(ii),  $\ell(R/I)=\ell(\widehat{R}/\mathcal{I})$. Now, we show $I$ is full parameter:
$$\mu(I)=\dim _{\frac{R}{\fm}}(\frac{I}{\fm I})=\dim ((\frac{I}{\fm I})\otimes \widehat{R})=\dim _{\frac{\widehat{R}}{\widehat{\fm}}}  (\frac{\mathcal{I}}{\widehat{\fm} \mathcal{I}})=\mu(\mathcal{I}).$$This implies that $\m(R)\leq\m(\widehat{R})$. Let $I\lhd R$ be a parameter ideal. Then $\widehat{I}\lhd \widehat{R}$ is a parameter ideal. Conclude from this that $\m(R)\geq\m(\widehat{R})$. So, $\m(R)=\m(\widehat{R})$.
\end{proof}

\begin{definition}Let $R$ be any local ring.
We say $R$ is homologically exact if $\D_h(R)=\m(R)$.
\end{definition}

The following may answer Question \ref{mat2}(ii):

\begin{proposition}
Let $R$ be a 1-dimensional local ring. Then $R$ is exact if and only if $R$ is homologically exact.
\end{proposition}

\begin{proof}Note that $\ell(R/xR)$ is finite if and only if $x$ is parameter. From this, $\m(R)=\rr(R)$.
By Corollary \ref{mc}, $\m(R)=\m(\widehat{R})$. It follows from Lemma \ref{comp} that  $\D_h(R)=\D_h(\widehat{R})$. Hence, without loss of the generality, we may assume that $R$ is complete.
By Fact \ref{G}
$\D(R)=\D_h(R)$.
It is now clear  that  $R$ is exact if and only if $R$ is homologically exact.
\end{proof}

\begin{lemma}\label{ab}
 Let $(R,\fm,k)$ be a local ring and  $M$  a module such that $\fm M=0$. Then $M$ is exact.
\end{lemma}

\begin{proof}
The condition $\fm M=0$ implies that  $M/xM=M$ for any $x\in\fm$. Hence, $\ell(M)=\rr(M).$
Also, $\fm M=0$ implies that
any
submodule $N$ of $M$ is  a vector space.
It turns out that $\dim_{k}N=\ell(N)=\ell(N/\fm N)=\mu(N)$.
Thus, $\D(M)=\dim_{k}M$. So,
$\D(M)=\dim_{k}M=\ell(M)=\rr(M).$
\end{proof}

\begin{lemma}\label{III}
Let $(R,\fm)$ be a generalized Cohen-Macaulay  ring of  dimension $d>0$ and that $k$ is an infinite. The following assertions hold: \begin{enumerate}
\item[i)] One has $\m(R)\leq\e(R)+\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R)).$
\item[ii)]  If  $R$ is Buchsbaum, then $\m(R)=\e(R)+\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R)).$
\end{enumerate}
\end{lemma}

\begin{proof}
i) For each parameter ideal $\fq$, set $I(\fq) := \ell(R/ \fq ) -\e(\fq,R)$. Since $R$ is generalized Cohen-Macaulay,
$I(R):=\sup\{I(\fq) \}$ is finite. In fact, $I(R)=\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R)).$
In view of \cite[Theorem 14.14]{Mat}, $\fm$ has a reduction $\fq_0$ which is   a parameter ideal.
Due to \cite[Theorem 14.13]{Mat} $\e(R)=\e(\fq_0,R)$. Putting all of these together, we have
\[\begin{array}{ll}
\m(R)&\leq\ell(R/ \fq_0 )\\
&=I(\fq_0)+\e(\fq_0,R)\\
&=I(\fq_0)+\e(R)\\
&\leq\sup\{I(\fq) \}+\e(R)\\
&=\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R))+\e(R).
\end{array}\]

ii)   Let $\fq_0$ be a parameter sequence such that $\m(R)=\ell(R/ \fq_0 )$.
 The natural surjection $R/\fq_0^n\to R/\fm^n$ shows that $\ell(R/\fq_0^n)\geq \ell(R/\fm^n)$.
Consequently, $\e(R)\leq\e(\fq_0,R)$. Also, $I(\fq_0)=\sup\{I(\fq):\fq\textit{ is parameter}\}$, because over Buchsbaum rings $I(\fq)$ is independent of $\fq$.  Putting these together,
\[\begin{array}{ll}\m(R)&=\ell(R/ \fq_0 )\\
&=I(\fq_0)+\e(\fq_0,R)\\
&\geq I(\fq_0)+\e(R)\\
&=\sup\{I(\fq) \}+\e(R)\\
&=\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R))+\e(R)\\
&\stackrel{(i)}\geq\m(R).
\end{array}\]
Thus, $\m(R)=\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R))+\e(R)$.
\end{proof}

The following is a higher-dimensional version of  \cite[Theorem 2.2]{ii}:

\begin{proposition}
Any  Buchsbaum ring with infinite residue field is homologically exact.
\end{proposition}

\begin{proof}
In view of \cite[Lemma 1.13]{bus}  a ring is Buchsbaum if and only its $\fm$-adic completion is Buchsbaum. By Corollary \ref{mc} $\m(R)=\m(\widehat{R})$. We may assume that $R$ is complete. Let $A$ be a Gorenstein ring of same dimension as of $R$ that maps onto $R$.
By  Lemma \ref{III}
$\m(R)=\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R))+\e(R).$ Since $R$ is Buchsbaum $\fm\HH^i_{\fm}(R)=0$ for al $i<\dim R$.
In view of Lemma \ref{ab} $\ell(\HH^i_{\fm}(R))=\mu(\HH^i_{\fm}(R))=\D(\HH^i_{\fm}(R))$. Since  $\dim (\HH^i_{\fm}(R))=0$, we have $\D(\HH^i_{\fm}(R))=\D_h(\HH^i_{\fm}(R))$.
In view of \cite[Corollary 4.2.3]{dil} Dilworth number of a finite length module is stable under taking Matlis duality.   In view of independence theorem and local duality from
Grothendieck's local cohomology modules, we have:
 \[\begin{array}{ll}
\m(R)&=\sum_{i=0}^{d-1}(^{d-1}_i)\ell(\HH^i_{\fm}(R))+\e(R)\\
&=\sum_{i=0}^{d-1}(^{d-1}_i)\D_h(\HH^i_{\fm}(R))+\e(R)\\
&=\sum_{i=0}^{d-1}(^{d-1}_i)\D_h(\HH^i_{\fm_A}(R))+\e(R)\\
&=\sum_{i=0}^{d-1}(^{d-1}_i)\D_h(\HH^i_{\fm_A}(R)^v)+\e(R)\\
&=\sum_{i=0}^{d-1}(^{d-1}_i)\D_h(\Ext^{d-i}_A(R,A))+\e(R).
\end{array}\]The last one is
$\D_h(R)$. This completes the proof.
\end{proof}
An example of homologically exact ring which is not Buchsbaum:
\begin{example}
Look at $R := \frac{\mathbb{Q}[[x, y]]}{x\fm^2}$. This is well known that
$\e(R)=\sup\{\mu(\fm^n):n\gg 0\}$. The set $\{x^iy^j:0<i+j=n\}$ is a generator for $\fm^n$. Due to $x^3=x^2y=xy^2=0$ we observe that $\fm^n=y^nR$ for all $n\geq3$. From this,
$\e(R)=1$. Also, $$
R/yR\simeq\frac{\mathbb{Q}[[x,y]]/(x^3,x^2y,xy^2)}{(y,x^3,x^2y,xy^2)/(x^3,x^2y,xy^2)}
\simeq \mathbb{Q}[[x,y]]/(y,x^3,x^2y,xy^2)
\simeq\mathbb{Q}[[x,y]]/(y,x^3).$$
As a vector-space, the last one is $\mathbb{Q}\oplus \mathbb{Q}x\oplus \mathbb{Q}x^2$. Thus, $\m(R)\leq\ell(R/yR)\leq3$. Note that $\HH^0_{\fm}(R)=xR$. This is principal. But its submodule $\fm\HH^0_{\fm}(R)=(x^2,xy)$ is not principal.
Consequently, $\D(\HH^0_{\fm}(R))>1$. Recall from \cite[Proposition 2.4]{ii} that $\D(R)=\e(R)+\D(\HH^0_{\fm}(R))$ and $\D(R)\leq\m(R)$. Putting these together: $3\leq\D(R)\leq\m(R)\leq3$.
Hence $\D(R)=\m(R)$. Therefore, $R$ is homologically exact. But $R$ is not Buchsbaum, because $\fm\HH^0_{\fm}(R)\neq 0$.
\end{example}
Rings of Dilworth number 2 are so special:

\begin{observation}(After Bass-Matlis)
Let $R$ be  a 1-dimensional reduced  ring  and
$R'$ be its normalization such that $\mu_R(R')<\infty$.  The following are equivalent:
a) $\D(R)=2$,
b) Any ring between $R$ and $R'$ is Gorenstein,
c) Any ring between $R$ and $R'$ is complete-intersection.
\end{observation}

\begin{proof}
The equivalence $a)\Leftrightarrow b)$ is in \cite[\S 7]{bass}.
Assume b). Let $A$ be any ring  between $R$ and $R'$.  Any ring between $A$ and $A'=R'$ is Gorenstein. By $b)\Leftrightarrow a)$ we have $\D(A)=2$.  Let $\fm$ be any maximal ideal of $A$. Then $\mu(\fm A_{\fm})\leq 2$. Reduced rings satisfy in the Serre's condition $R(0)$. If $\mu(\fm A_{\fm})\leq 1$ then $A_{\fm}$ is regular. Without loss of the generality  we may assume that $\mu(\fm A_{\fm})=2$.
Note $A_{\fm}$ is reduced. Thus, $\dim A_{\fm}\neq 0$. In particular,   $A_{\fm}$ is  Cohen-Macaulay and that $\emb(A_{\fm})=\dim A_{\fm}+1$. In view of \cite[Ex. 21.2]{Mat}, $A_{\fm}$ is complete-intersection. In any cases $c)$ follows.
Clearly, $c)$ implies b).
\end{proof}

The reduced assumption is important:
\begin{example}
Look at $R := \frac{\mathbb{Q}[[x, y]]}{(x^2, xy)}$. Then $\depth (R)=0$ and $\dim R=1$.
Note that  $\e(R)=\lim_{n\to\infty}
\frac{\ell(R/ \fm^n)(d - 1)^!}{n^{d-1}}=1.$
 One has $\rad(y)=(x,y)$. Thus $y$ is a parameter element. Also,$$ R/yR=\frac{\mathbb{Q}[[x,y]]/(x^2,xy)}{(x^2,y)/(x^2,xy)}\simeq \mathbb{Q}[[x,y]]/(x^2,y)\simeq_{vector-space}\mathbb{Q}\oplus \mathbb{Q}x.$$Thus, $\ell(R/yR)\leq2$. Hence, $1\leq\m(R)\leq2$. However, $\m(R)\neq1$ because $R$  is not regular. Therefore, $\m(R)=2.$
 On the other hand $\D(R)\leq\m(R)=2$ and that $\mu(\fm)=2$.  So, $\D(R)=2$. But, $R$ is not even Cohen-Macaulay. So, the implication $a)\Rightarrow b)$
 in the above observation is not true without the reduced assumption.
\end{example}

Recall that a system of parameters for a non-zero finitely generated
$R$-module $M$ is a sequence $\underline{x}:=x_1,\ldots,x_n$  where $n:=\dim M$   such that
$M/\underline{x}M$ is artinian. By $\m(M)$ we mean $\inf\{\ell(M/\underline{x}M):\underline{x}\textit{ is a  M-parameter sequence} \}$ and we call it  \textit{Matsumura number} of $M$.
Also, $M$ is called homologically exact if $\m(M)=\D_h(M)$.

\begin{lemma}
Let $R$ be a local ring and $M$ an module of dimension same as of $\dim R$. Then
$M$ is maximal Cohen-Macaulay if and only if  $\e(M)=\D_h(M)$.
\end{lemma}

\begin{proof}
We may assume $R$ is complete. Let $d:=\dim M=\dim R$. Suppose $R$ is  homomorphic image of a Gorenstein ring $A$ of same dimension  as of $R$.
 First, assume that $\e(M)=\D_h(M)$.
 Then $\Ext^{d-i}_A(M,A) =0$ for all $i<d$.
By local duality, $\HH^i_{\fm_A} (M)=0$ for all $i<d$. By the independence theorem, $\depth_R(M)=d$. So, $M$ is maximal Cohen-Macaulay.
Conversely, assume that $M$ is maximal Cohen-Macaulay. Let $\underline{x}$ be a parameter sequence for $R$. Let $\pi:A\to R$ be the surjection. There are $y_i$ such that $\pi(y_i)=x_i$.
Note that $\rad(y_i)=\fm_A$. Hence, $\underline{y}:=\{y_i\}$ is a parameter sequence for $A$, because $\dim A=\dim R$. We view  $M$ as an $A$-module via $\pi$. By this, $\underline{y}$ is $M$-sequence, i.e.,
$M$ is maximal Cohen-Macaulay as an $A$-module. Then $\Ext^i_A(M,A)=0$ for all $i>0$. By definition,  $\e(M)=\D_h(M)$.
\end{proof}

The following is a higher-dimensional version of \cite[Theorem 3.3]{ii}:

\begin{proposition}
Let $(R,\fm,k)$ be a local Cohen-Macaulay ring with infinite residue field and with a canonical module $\omega_R$. Then
$$\m(\omega_R)=\D_h(\omega_R)=\e(\omega_R)=\m(R)=\D_h(R)=\e(R).$$
\end{proposition}

\begin{proof}
Since $\omega_R$ is maximal Cohen-Macaulay and by  the above lemma,
$\e(\omega_R)=\D_h(\omega_R)$.
Let $\fq$ be any $\fm$-primary ideal. Recall that the  natural surjection $\frac{\omega_R}{\fq^n\omega_R}\to \frac{\omega_R}{\fm^n\omega_R}$ shows that $\e(\fq,\omega_R)\geq \e(\omega_R)\ \ (\ast)$.
 Let $\underline{x}$ be a system of parameters.   By a theorem of Serre, $\Sigma_{i>0}(-1)^{i+1}\ell(\HH^i(\underline{x},\omega_R))\geq 0$, i.e., $\e(\underline{x},\omega_R)\leq \ell(\omega_R/\underline{x}\omega_R)$. Let $\underline{x}_0$ be such that the $\inf$ in $\m(\omega_R)$ happens.
Thus, $$\m(\omega_R)=\ell(\omega_R/\underline{x}_0\omega_R)\geq\e(\underline{x}_0,\omega_R)\stackrel{(\ast)}\geq\e(\omega_R).$$Since $k$ is infinite,
 there is a reduction of $\fm$ by a parameter sequence $\underline{x}$ (see \cite[Theorem 14.14]{Mat}).
In view of \cite[Theorem 14.13]{Mat}, $\e(\omega_R)=\e(\underline{x},\omega_R)$. Since $\underline{x}$ is $\omega_R$-sequence,
$\Sigma_{i>0}(-1)^i\ell(\HH^i(\underline{x},\omega_R))= 0$, i.e.,
 $\e(\underline{x},\omega_R)=\ell(\omega_R/\underline{x}\omega_R)$. Hence  $$\e(\omega_R)=\e(\underline{x},\omega_R)=\ell(\omega_R/\underline{x}\omega_R)\geq\m(\omega_R).$$
Therefore,  $\m(\omega_R)=\D_h(\omega_R)=\e(\omega_R)$.
Similarly, $\m(R)=\D_h(R)=\e(R)$. If $\omega_R$ has a rank then its rank is one and so $\e(\omega_R)=\e(R)$. But $\omega_R$ has a rank if and only if $R$ is generically Gorenstein.
Let us handle the general case by the following trick of Ikeda. Let $\underline{x}$ be any parameter sequence.  Then  $\underline{x}$ is  both $R$-regular and $\omega_R$-regular. Also, $\frac{\omega_R}{\underline{x}\omega_R}\simeq\omega_{R/\underline{x}R}$. Recall that $\overline{R}:=R/\underline{x}R$ is zero-dimensional. In particular,  it is complete and the only  indecomposable injective module is $E_{\overline{R}}(k)$. Thus, $\omega_{\overline{R}}\simeq E_{\overline{R}}(k)$. In view of Matlis duality $$\ell(\overline{R})=\ell(\overline{R}^v)=\ell(E_{\overline{R}}(k))=\ell(\omega_{\overline{R}})=\ell(\frac{\omega_R}{\underline{x}\omega_R})\quad(+)$$

Let $\underline{y}=y_1,\ldots,y_d\subset\fm$ be any sequence of elements. If
 $\underline{y}$
is an $R$-parameter sequence,  then $\ell(R/\underline{y}R)<\infty$. From this $\ell(R/\underline{y}R\otimes\omega_R)<\infty$. Since for any
strict subset $\underline{y}'$ of  $\underline{y}$ we have $\ell(R/\underline{y}'R\otimes\omega_R)=\infty$, it is clear that $\underline{y}$
is an $\omega_R$-parameter sequence. Conversely, suppose $\underline{y}$
is an $\omega_R$-parameter sequence. Then $\underline{y}$ is an $\omega_R$-regular sequence. (If $\pd(\omega_R)<\infty$, then by  Auslander zero-divisor, $\underline{y}$ is an $R$-regular sequence. But, $\pd(\omega_R)<\infty$ if and only if $R$ is Gorenstein. Instead, we use the following trick:) Recall that
 $\frac{\omega_R}{(y_1,\ldots,y_i)\omega_R}$ is Cohen-Macaulay for all $i$.
For any $\fp\in\Ass(\frac{\omega_R}{(y_1,\ldots,y_i)\omega_R})$ we have $\dim R/ \fp=d-i$.
Note that canonical module has  support same as of the prime spectrum. Also, associated prime
ideals of a Cohen-Macaulay module is the minimal elements of its supports. From these,  $$\Ass(\frac{\omega_R}{(y_1,\ldots,y_i)\omega_R})=\Ass(\frac{R}{(y_1,\ldots,y_i)}).$$
This yields that $\dim (\frac{R}{(y_1,\ldots,y_i)})=d-i$ for all $i$.
It turns out that  $\underline{y}$
is an $R$-parameter sequence. 

Combine the previous  paragraph along with $(+)$
to conclude that $\m(\omega_R)=\m(R)$. Consequently: $$\m(\omega_R)=\D_h(\omega_R)=\e(\omega_R)=\m(R)=\D_h(R)=\e(R).$$
The proof is now complete.\end{proof}

\begin{remark}
i) Here we present a (new) proof of Serre's positivity theorem: $\Sigma_{i>0}(-1)^{i+1}\ell(\HH^i(\underline{x},R))\geq 0$.
Indeed,
recall that $I:=(x_1,\ldots,x_d)$ is a full parameter ideal. We may assume that $d>0$. The assignment $X_i\mapsto x_i + (\underline{x})^2$ gives the surjection
$P:= \frac{R}{I}[X_1, \ldots ,X_d]\twoheadrightarrow \gr_{I}(R) =: G$.
 This yields that
$$\ell(R/I)
= \lim_{n\to\infty} \frac{\ell(P_n)(d - 1)^!}{n^{d-1}}
\stackrel{(\ast)}\geq \lim_{n\to\infty}
\frac{\ell(G_n)(d - 1)^!}{n^{d-1}}=\lim_{n\to\infty}
\frac{\ell(R/I^n)(d - 1)^!}{n^{d-1}}=\e(I,R).$$
Since $\e(I,R)=\Sigma_{i\geq0}(-1)^i\ell(\HH^i(\underline{x},R))$, we see $\Sigma_{i>0}(-1)^i\ell(\HH^i(\underline{x},R))\leq 0$. So, the claim follows.

ii) Recall that $(\ast)$ in part i) is the equality if and only if $\underline{x}$ is a  regular sequence. From this we can reprove the following result of Shoutens:
Let $(R,\fm)$ be  a local ring  such that $|R/\fm|=\infty$. Then $R$ is Cohen-Macaulay
if and only if $\m(R)=\e(R)$.
\end{remark}

\section{Concluding remarks and questions  }

Are zero-dimensional complete-intersections of zero-characteristic exact?

\begin{fact}(Watanabe  et al.)
	Let $(R,\fm,k)$ be an artinian equi-characteristic local ring.
	\begin{enumerate}
		\item[i)] Any homogeneous  complete-intersection of characteristic zero and $\emb(R)=3$ is exact.
		\item[ii)] If $f_1,\ldots,f_n\in \mathbb{Q}[X_1,\ldots,X_n]$ be homogeneous  regular sequence and of degree $d_1,\ldots,d_n$ with
$d_1\leq\ldots\leq d_n$ with $d_n\geq d_1+\ldots d_{n-1}-n$. Then $R:=\frac{\mathbb{Q}[X_1,\ldots,X_n]}{(f_1,\ldots ,f_n)}$ is exact. 	
		\item[iii)]
 If $d_n\geq d_1+\ldots d_{n-1}-n+2$
the same claim as of ii)  holds for any infinite field.\item[iv)] If $f_1,\ldots,f_n\in \mathbb{Q}[X_1,\ldots,X_n]$ is a monomial complete-intersection,
then $R:=\frac{\mathbb{Q}[X_1,\ldots,X_n]}{(f_1,\ldots ,f_n)}$ is exact.
\end{enumerate}
\end{fact}

\begin{proof}	i)  The ring $R$ is of the form $\frac{K[x,y,z]}{(f_1,f_2,f_3)}$ where  $f_i$ are homogeneous and of degree $d_i$. We may assume that $2\leq d_1\leq d_2\leq d_3$. By the proof of \cite[Theorem 3.48]{wat3}, $\rr(R)=\sup\{\mu(\fm^i)\}$. That is $R$ is exact.

ii) By the induction assumption we may assume  that $2\leq d_1$. By the proof of \cite[Corollary 3.54]{wat3}, $\rr(R)=\sup\{\mu(\fm^i)\}$. 

iii) This is in \cite[Remark 3.55]{wat3}.

iv) Monomial complete-intersections over fields of zero-characteristic has Lefschetz property.
By \cite[Proposition 3.6]{wat3} $\rr(R)=\sup\{\mu(\fm^i)\}$. So, $R$ is exact.
\end{proof}

It may be $\D(R)\neq\sup\{\mu(\fm^n)\}$ but $R$ is exact. Such a thing happens even over Gorenstein rings. The first example was found by Ikeda \cite[Example 4.4]{i}.
Also, Ikeda conjectured  that $\D(R)=\sup\{\mu(\fm^n)\}$ over any artinian Gorenstein ring of embedding dimension 3. There are
counter-examples in the positive characteristic, see \cite{bcan}.
Let $(R,\fm)$ be a zero-dimensional Gorenstein local ring and of embedding dimension $3$ over a field of zero-characteristic. Is $\D(R)=\sup\{\mu(\fm^n)\}$?

\begin{remark}
i) 	Let $\mathbb{Q}\subset R$ be an artinian Gorenstein local ring such that $\mu(\fm)=\mu(\fm^2)=3$ and that  $\fm^4=0$. Then $R$ is exact.	In fact $\D(R)=\sup\{\mu(\fm^n)\}$.
Indeed, we may assume that $\fm^3$ is not zero. Also, $\mu(\fm^3)=1$. In particular,
 $\max\{\mu(\fm^i) \}=3$. By the proof of Corollary \ref{8} we have $\D(R)=\rr(R)=3$.

ii) The assumption $\mathbb{Q}\subset R$ in part i) is important.
	Indeed, let  $R:=\frac{\mathbb{F}_2[X,Y,Z]}{(X^2,Y^2,Z^2)}$.  Clearly, $R$ is an artinian Gorenstein local ring such that $\mu(\fm)=\mu(\fm^2)=3$ and that  $\fm^4=0$. We recall from \S 2 that $R$ is not exact.
 \end{remark}

By $\hdeg(-)$ we mean the homological degree. 
When is
$\m(R)\leq\hdeg(R)$?

\begin{remark}Let $(R,\fm,k)$ be an analytically irreducible local ring with infinite residue field. Then
$\m(R)\leq\hdeg(R)$.
\end{remark}

\begin{proof}If $R$ is of equi-characteristic the claim is in \cite{sch}. So, we may assume that $\Char R=0\neq p=\Char k$. We apply the same argument as of \cite{sch}.
Since both of $\m(R)$ and $\hdeg(R)$ behave nicely with respect to the completion, we may assume that $R$ is a complete local domain.
Let  $\underline{x}:=p, y_2,\ldots,y_d$ be a parameter sequence for $R$. By Cohen's structure theorem, there is a regular local ring $(A,\fn)\subset R$
such that  $\underline{x}A=\fn$ and that $R$ is finitely generated as a module over $A$. In fact, $A=V[[y_2,\ldots,y_d]]$ where $V$ is the coefficient ring of $R$ which is a discreet  valuation ring.  It follows
from  \cite[Proposition 4.1]{vas} that $\mu_A(R)\leq\hdeg(R)$. In sum, $\m(R)\leq \ell(R/\underline{x}R)=\ell(R/ \fn R)=\mu_A(R)\leq\hdeg(R),$ as claimed.
\end{proof}

\begin{acknowledgement}
I thank
Junzo Watanabe for his valuable comments on the rough version of this draft.

\end{acknowledgement}

\end{document}